a can has 3tennis balls, the can is just large enough sothat he tenis balls will fit inside with the lid on. the diameter of each tennis ball is 2.5 inches. what is te volume of the empty space inside the can?
Answer by mananth(16946)

Volume of cylinder - volume of three balls = volume of empty space.
Cylinder = pi*r^2*h
The height of cylinder = 3 times the diameter of the ball.
The radius of cylinder = radius of ball
= 3.14 * 1.25^2*7.5 =36.8 cu.in

Volume of 1 ball = 4/ 3 * pi* *r^3
=4/3 * pi* 1.25^3
=8.18 cu.in
3 balls volume = 8.18 *3
= 24.54 cu.in
empty space = 36.8 - 8.18= 27.68 cu.in